Consider a hollow spherical conductive shell of radius (R) 0.2 m with a fixed



charge of +2.0 x 10-6 C uniformly distributed on its surface.



a. What is the electric field at all points inside the sphere? Express your answer as



a function of the distance r from the center of the sphere. (2 points)



b. What is the electric field outside the sphere? Express your answer as a function



of the distance r from the center of the sphere. (2 points)



c. What if the sphere is a solid conductive sphere? What is the electric field at all



points inside the sphere? Express your answer as function of the distance r from



the center of the sphere

1
Expert's answer
2022-06-06T08:53:03-0400

(a) Inside hollow conducting sphere the electric field is zero at all points.

(b) Outside sphere the electric field is given by

"E=k\frac{q}{r^2}\\
=9*10^9*\frac{2.0*10^{-6}}{r^2}=\frac{18000}{r^2}\:\rm \frac{V}{m}"

(c) Inside solid conducting sphere the electric field is zero at all points.